A local FedEx Office store has a press that can print 5,800 brochures per hour. How many can be printed during a 3 1/4 -hour run
Answer by stanbon(75887)

A local FedEx Office store has a press that can print 5,800 brochures per hour. How many can be printed during a 3 1/4 -hour run
-----
Use a proportion::
#/time = #/time
-----
x/[3 1/4] = 5800/1
-----
x = (3 1/4)(5800) = (13/4)(5800) = 13*1450 = 18,850
=======================
Cheers,
Stan H.
=======================